# Influence Functions for Edge Edits in Non-Convex Graph Neural Networks

### v1

@@ Appendix B Gradient derivation for edge-edit PBRF
Similar to the derivation of the influence function for the PBRF objective in Section A.2, we demonstrate that the influence function in Equation 10 corresponds to that of the linearized form of the edge-edit PBRF objective:
@@
which confirms that the linearized edge-edit PBRF objective yields the same influence function as derived in Equation 10.
Appendix C Description of LiSSA
@@
Appendix D Experimental result on other datasets and GNNs
In this section, we present scatter plots for additional non-convex GNNs and datasets. Figure 6 shows the results for additional datasets, while Figure 7 and Figure 8 show the results for ChebNet [12] and GAT [31], respectively. Under these settings, our influence function accurately predicts the actual influence, consistently showing a correlation above $0.8$.
@@ Appendix D Experimental result on other datasets and GNNs
Appendix E Analysis on FoSR
In this section, we analyze the edge rewiring strategy of FoSR [18] using our influence function. FoSR rewires the graph by inserting edges that increase the spectral gap. Figure 9 presents the influence of the inserted edges in terms of validation loss, over-squashing, and Dirichlet energy.
(a) FoSR [18]
Figure 9: Histograms of the estimated influence of edge insertions selected by FoSR [18], evaluated on a four-layer GCN trained on the Texas dataset. Influences are measured across validation loss, over-squashing, and over-smoothing. Arrows ($\downarrow$/$\uparrow$ / ↑) indicate the desired direction of each measurement (decrease/increase).
Among the 200 edges inserted by FoSR, 143 are predicted to reduce over-squashing, suggesting that FoSR effectively identifies and adds edges that mitigate this issue. However, a similar limitation to that observed in BORF [24] emerges: the inserted edges do not necessarily improve the validation loss and may even exacerbate over-smoothing. These results consistently show that mitigating over-squashing does not necessarily improve other evaluation metrics, highlighting the necessity of considering multiple perspectives.
@@ Appendix F Experimental configuration
All experiments are conducted using NVIDIA GeForce RTX 3090, NVIDIA RTX A5000, and NVIDIA RTX A6000 GPUs. The experiments presented in the main text employ a 4-layer GCN model as a representative non-convex GNN. To produce the results in Table 1 and Table 2, we tune the model and training hyperparameters of a vanilla GCN over the following search space: learning rates $\{0.1,0.03,0.01\}$, hidden dimensions $\{32,64\}$, and weight decays $\{10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\}$. Training is performed for 2000 epochs using the SGD optimizer. For influence function computation, we run the LiSSA algorithm for 10,000 iterations. The damping parameter $\lambda$ is selected from $\{0.1,0.01,0.001,0.0001\}$. We randomly sample 10,000 candidate edges for both deletion and insertion, and estimate their influence. For Table 1, the number of edges to edit is determined based on validation accuracy. The experiments are repeated for 10 independent runs using different random seeds provided by the BernNet implementation.
BORF [24] and FoSR [18], used in the analysis of edge rewiring methods, are applied with the default settings from the original paper. The number of rewired edges reported in Figure 4 is aggregated over 10 runs, with a total of 200 edges inserted. Unlike FoSR, which performs only edge insertions, BORF also considers edge deletions. To improve visualization, we reduce the number of edge deletions, as the influence scores for insertions exhibit a long-tailed distribution, making it difficult to display them on the same scale as deletions. Accordingly, edge deletions are performed with 120, 100, and 100 edges, respectively.
For the adversarial attack experiments in Table 2, DICE [32] and PRBCD [13] are implemented by modifying the PyTorch-based DeepRobust library [22], while maintaining its default settings. For our method, we consider both edge insertions and deletions, and select the operations with the highest influence scores for validation loss, and the lowest influence scores for over-squashing and Dirichlet energy.

### v2

@@ Appendix B Gradient derivation for edge-edit PBRF
Similar to the derivation of the influence function for the PBRF objective in Section˜A.2, we demonstrate that the influence function in Equation˜10 corresponds to that of the linearized form of the edge-edit PBRF objective:
@@
which confirms that the linearized edge-edit PBRF objective yields the same influence function as derived in Equation˜10.
Appendix C Influence function for multiple edge edits
Problem setup for multiple edge edits
In this section, we extend our influence function to handle multiple edge edits. Let $\mathcal{S}$ denote the set of edges to be edited. We generalize the definition of $A^{\epsilon}$ in Section˜3 to the multi-edge case by defining
$A^{\epsilon}_{uv}=A_{uv}+\bigl(2\mathbb{I}[\{u,v\}\in\mathcal{S}]-1\bigr)N\epsilon.$
The edge-reweighted graph is then given by $\mathcal{G}^{\epsilon}=\{\mathcal{V},\mathcal{E},A^{\epsilon}\}$, where edges in $\mathcal{S}$ are reweighted. Setting $\epsilon=-1/N$ implies that all edges in $\mathcal{S}$ are deleted if they exist, and inserted otherwise.
Parameter shift
Following Section˜3, the edge-edit PBRF is defined as
Following the same derivation steps as in Appendix˜B, the contribution of the parameter shift to the evaluation function is given by
$\nabla_{\theta}f(\theta^{*}_{0},\mathcal{G})^{\top}\left.\frac{\partial\theta^{*}_{\epsilon}{}}{\partial\epsilon}\right|_{\epsilon=0}=-\nabla_{\theta}f(\theta_{s},\mathcal{G})^{\top}\mathbf{G}^{-1}\sum_{v\in\mathcal{V}_{\text{train}}}\left(\nabla_{\theta}\mathcal{L}\left(h_{v}^{\mathcal{G},\theta_{s}}\right)-\nabla_{\theta}\mathcal{L}\left(h_{v}^{\mathcal{G}^{-\frac{1}{N}},\theta_{s}}\right)\right),$ (34)
where $\mathbf{G}=\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}+\lambda\mathbf{I}$. The only distinction from the single-edge case lies in the redefinition of $\mathcal{G}^{-1/N}$, which now denotes the graph with multiple edge edits.
Message propagation
Unlike the single-edge case in Section˜3, the perturbed adjacency matrix $A^{\epsilon}$ is influenced simultaneously by all edges in the set $\mathcal{S}$. Consequently, the message propagation effect can no longer be attributed to a single edge but must be decomposed into contributions from all edited edges. By applying the chain rule, the overall effect is expressed as a summation over gradients with respect to each reweighted edge:
$\left.\frac{\partial f(\theta,\mathcal{G}^{\epsilon})}{\partial A^{\epsilon}{}}\frac{\partial A^{\epsilon}{}}{\partial\epsilon}\right|_{\theta=\theta^{*}_{0},\;\epsilon=0}$ $=\sum_{\{u,v\}\in\mathcal{S}}\left.\frac{\partial f(\theta,\mathcal{G}^{\epsilon})}{\partial A^{\epsilon}_{uv}{}}\frac{\partial A^{\epsilon}_{uv}{}}{\partial\epsilon}\right|_{\theta=\theta_{s},\,\epsilon=0}+\left.\frac{\partial f(\theta,\mathcal{G}^{\epsilon})}{\partial A^{\epsilon}_{vu}{}}\frac{\partial A^{\epsilon}_{vu}{}}{\partial\epsilon}\right|_{\theta=\theta_{s},\,\epsilon=0}$
$=\sum_{\{u,v\}\in\mathcal{S}}(2\mathbb{I}[\{u,v\}\in\mathcal{E}]-1)\,N\left(\frac{\partial f(\theta_{s},\mathcal{G})}{\partial A_{uv}}+\frac{\partial f(\theta_{s},\mathcal{G})}{\partial A_{vu}}\right).$ (35)
Unified influence function under multiple edge edits
By substituting the parameter shift and message propagation terms into Equation˜8 and linearizing around $\epsilon=0$, the resulting influence under multiple edge edits can be approximated as
$\quad-\sum_{\{u,v\}\in\mathcal{S}}(2\mathbb{I}[\{u,v\}\in\mathcal{E}]-1)\left(\frac{\partial f(\theta_{s},\mathcal{G})}{\partial A_{uv}}+\frac{\partial f(\theta_{s},\mathcal{G})}{\partial A_{vu}}\right).$ (36)
Appendix D Description of LiSSA
@@
Appendix E Experimental result on other datasets and GNNs
In this section, we present scatter plots for additional non-convex GNNs and datasets. Figure˜7 shows the results for additional datasets, while Figure˜8 and Figure˜9 show the results for ChebNet [13] and GAT [33], respectively. Under these settings, our influence function accurately predicts the actual influence, consistently showing a correlation above $0.8$.
@@ Appendix F Experimental configuration
All experiments are conducted using NVIDIA GeForce RTX 3090, NVIDIA RTX A5000, and NVIDIA RTX A6000 GPUs. The experiments presented in the main text employ a 4-layer GCN model as a representative non-convex GNN. To produce the results in Table˜1 and Table˜2, we tune the model and training hyperparameters of a vanilla GCN over the following search space: learning rates $\{0.1,0.03,0.01\}$, hidden dimensions $\{32,64\}$, and weight decays $\{10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\}$. Training is performed for 2000 epochs using the SGD optimizer. For influence function computation, we run the LiSSA algorithm for 10,000 iterations. The damping parameter $\lambda$ is selected from $\{0.1,0.01,0.001,0.0001\}$. We randomly sample 10,000 candidate edges for both deletion and insertion, and estimate their influence. For Table˜1, the number of edges to edit is determined based on validation accuracy. The experiments are repeated for 10 independent runs using different random seeds provided by the BernNet implementation.
BORF [26] and FoSR [19], used in the analysis of edge rewiring methods, are applied with the default settings from the original paper. The number of rewired edges reported in Figure˜6 is aggregated over 10 runs, with a total of 200 edges inserted. Unlike FoSR, which performs only edge insertions, BORF also considers edge deletions. To improve visualization, we reduce the number of edge deletions, as the influence scores for insertions exhibit a long-tailed distribution, making it difficult to display them on the same scale as deletions. Accordingly, edge deletions are performed with 120, 100, and 100 edges, respectively.
For the adversarial attack experiments in Table˜2, DICE [34] and PRBCD [14] are implemented by modifying the PyTorch-based DeepRobust library [24], while maintaining its default settings. For our method, we consider both edge insertions and deletions, and select the operations with the highest influence scores for validation loss, and the lowest influence scores for over-squashing and Dirichlet energy.